# Asymptotic Distribution-free Change-point Detection for Modern Data Based on a New Ranking Scheme

### v3

@@ IV-A The choice of k
The choice of graphs remains an open question for CPD based on similarity graphs [28, 30, 49]. We adapt the method in [49] and [48]. Specifically, they compare the empirical power of the method for different choices of $k=[n^{\lambda}]$ by varying $\lambda$ from 00 to $1$. [49] suggested to use $k=[n^{0.5}]$ for the generalized edge-count test (GET) [30] when the $k$-MST is used, while [48] recommended $k=[n^{0.65}]$ for their two-sample test statistic with graph-induced rank on the $k$-NNG. We follow the same way in choosing $k$ for $\rm M_{g}$-NN. We generate independent sequences from three different distribution pairs of $(F_{0},F_{1})$:
@@
    the multivariate Gaussian distribution $\big{(}N_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),N_{d}(\frac{30}{\sqrt{nd}}\mathbf{1}_{d},\mathbf{I}_{d})\big{)}$;
@@
    the multivariate $t_{3}$ distribution $\big{(}t_{3}(\mathbf{0}_{d},\mathbf{I}_{d}),t_{3}(\frac{30}{\sqrt{nd}}\mathbf{1}_{d},(1+\frac{30}{\sqrt{nd}})^{2}\mathbf{I}_{d})\big{)}$;
@@
    the multivariate Cauchy distribution $\big{(}{\rm Cauchy}_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),{\rm Cauchy}_{d}(\frac{30}{\sqrt{n}}\mathbf{1}_{d},\mathbf{I}_{d})\big{)}$.
The parameters are chosen so that the tests have moderate power. The change-point is set to be $\tau=n/2$, the dimension $d=500$ and $n=50,100,200$. We set $n_{0}=\lceil 0.05n\rceil$ and $n_{1}=n-n_{0}$, which will also be our choice by default in the latter experiments, where $\lceil x\rceil$ denotes the smallest integer larger than or equal to $x$. For comparison, we also include two graph-based methods, GET and the max-type edge-count test (MET) proposed in [30] using the $k$-MST. The empirical power is defined as the ratio of successful detection where the $p$-value is smaller than $0.05$. For fairness, the $p$-values are approximated by $1,000$ random permutations for all methods.
Figure 6 shows the power of these tests for $k=[n^{\lambda}]$, $\lambda\in(0,0.8]$. We see that the power of these tests first increases quickly when $k$ or $\lambda$ increases. If $k$ continues to increase, the power of GET and MET decreases dramatically, but the performance of $\rm M_{g}$-NN seems more robust. The reason may be that a denser graph can contain more similarity information, while noisier information can also be incorporated when more edges are included. However, $\rm M_{g}$-NN alleviates the problem by incorporating ranks on edges as those later induced edges have smaller ranks. The overall performance of $\rm M_{g}$-NN is the best, with a significant improvement of the power over heavy-tailed settings (the multivariate $t_{3}$ and Cauchy distributions) and the robustness over a wide range choice of $k$. Finally, we choose $\lambda=0.65$ for $\rm M_{g}$-NN, and $\lambda=0.5$ for GET and MET in the following analysis, which is reasonable for these methods to achieve adequate power and coincides with previous choices [49, 48].
@@ IV-B Performance comparison
We compare RING-CPD to GET and MET on $k$-MST with $k=[n^{0.5}]$, the graph-based method on the shortest Hamiltonian path [29] (SWR), the method based on Fréchet means and variances [59] (DM). We also compare three interpoint distance-based methods, the widely used distance-based method E-Divisive (ED) [26] implemented in the R package ecp, and the other two methods proposed recently by [27] and [34]. [27] proposed four statistics and we compare the statistic $C_{2N}$ that had a satisfactory performance in most of their simulation settings. [34] proposed three statistics, which perform well for location change, scale change, and general change, respectively. Here we compare with their statistic $S_{3}$, which they concluded to have relatively robust performance across various alternatives. For fairness, the $p$-values of all these methods are approximated by $1,000$ random permutations.
We set $n=200$ and the change-point $\tau=[n/3]$ and consider the dimension of the distributions $d=200,500,1000$. Before the change-point $y_{i}\sim F_{0}$ and after the change-point $y_{i}\sim F_{1}$. We consider both the empirical power and the detection accuracy estimated from $1000$ trials for each scenario. The empirical power is the ratio of the successful detection defined as $p$-value smaller than the nominal level $0.05$. The detection accuracy is provided in parentheses, which is the ratio of trials that the detected change-point is located in $[\tau-0.05n,\tau+0.05n]$ and the $p$-value smaller than $0.05$. We consider various settings that cover light-tailed, heavy-tailed, skewed, and mixture distributions for location, scale, and mixed alternatives. Specifically, we consider six settings for $F_{i}$, $i=0,1$:
@@
    the multivariate Gaussian distribution $N_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})$;
@@
    the multivariate Cauchy distribution ${\rm Cauchy}(\bm{\mu}_{i},\bm{\Sigma}_{i})$;
@@
    the multivariate $\chi_{5}^{2}$ distribution $\chi_{5}^{2}(\bm{\mu}_{i},\bm{\Sigma}_{i})$ (generated as $\bm{\Sigma}_{i}^{\frac{1}{2}}(X-5\mathbf{1}_{d}+\bm{\mu}_{i})$ where the $d$ components of $X$ are i.i.d. $\chi_{5}^{2}$);
@@
    the Gaussian mixture distribution $WN_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})+(1-W)N_{d}(-\bm{\mu}_{i},\bm{\Sigma}_{i})$ with $W\sim{\rm Bernoulli}(0.5)$;
@@
    the multivariate normal distribution with $t_{7}$ outliers $WN_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})+(1-W)t_{7}\big{(}\bm{\mu}_{i},\bm{\Sigma}_{i})$ with $W\sim{\rm Bernoulli}(0.9)$.
@@ V-A Seizure detection from functional connectivity networks
We illustrate RING-CPD for identifying epileptic seizures, which over two million Americans are suffering from [60]. As a promising therapy, responsive neurostimulation requires automated algorithms to detect seizures as early as possible. Besides, to identify seizures, physicians have to review abundant electro-encephalogram (EEG) recordings, which in some patients may be quite subtle. Hence, it is important to develop methods with low false positive and false negative rates to detect seizures from the EEG recordings. We use the “Detect seizures in intracranial EEG (iEEG) recordings” database by the UPenn and Mayo Clinic (https://www.kaggle.com/c/seizure-detection), which consists of the EEG recordings of $12$ subjects (eight patients and four dogs). For each subject, both the normal brain activity and the seizure activity are recorded multiple times, which are one-second clips with various channels (from $16$ to $72$), reducing to a multivariate stream of iEEGs. Following the preprocessing procedure of [9], we represent the iEEG data as functional connectivity networks using Pearson correlation in the high-gamma band ($70$-$100$Hz) [61]. Functional connectivity networks are weighted graphs, where the vertexes are the electrodes, and the weights of edges correspond to the coupling strength of the vertexes. An illustration of the networks is in Figure 7. The sample sizes of the $12$ subjects are also different, and the true change-point $\tau$’s are also known - before the change-point, the networks are from the seizure period, while after the change-point, the networks are from the normal brain activity, so we have the ground truth. We use the Frobenius norm to measure the distance between the observations represented by the weighted adjacency graphs.
@@
We do not include SWR in the comparison here since SWR does not perform well in the simulation studies and is time-consuming. For $C_{2N}$, since it does well under some simulation settings, we try to include it in the comparison. However, it is not only time-consuming but also memory-consuming (e.g., it requires at least $17$Gb size of memory when $n=1320$); we are only able to run it for $n\leq 600$, thus only showing its result for Dog $1$, and Patients $1$ and $4$. Since the sample size of each subject is large enough, we use the asymptotic $p$-value approximation for $\rm M_{g}$-NN and MET. We omit the result of GET since it performs similarly to MET but its $p$-value approximation is not as exact as MET. For DM, ED, and $S_{3}$, we still use $1,000$ random permutations to obtain the $p$-values. The results are summarized in Table V, where the absolute difference between the true change-point and the detected change-point $|\hat{\tau}-\tau|$ is reported. The $p$-values are not reported as they are smaller than $0.01$ for all methods and subjects. Our method achieves the same detection error as MET, which is very small for all subjects. ED also performs well, but with a slightly large error for Patient $4$. Although DM and $S_{3}$ achieve small errors for most subjects, they attain large detection errors for Patients $3$ and $4$. The performance of $C_{2N}$ is not robust in that it shows a large detection error for Patient $4$.
TABLE V: The absolute difference between the true change-point and the detected change-point ($|\hat{\tau}-\tau|$). The $p$-values of all methods for all subjects are smaller than $0.01$. $\rm M_{g}$-NN is the proposed method.
Subject $n$ $\tau$ $\rm M_{g}$-NN MET DM ED $C_{2N}$ $S_{3}$
@@
Patient 1 174 70 1 1 1 0 7 1
Patient 2 3141 151 7 7 13 1 - 13
Patient 3 1041 327 0 0 162 1 - 162
Patient 4 210 20 0 0 67 11 137 67
Patient 5 2745 135 3 3 5 2 - 5
Patient 6 2997 225 0 0 0 1 - 0
Patient 7 3521 282 2 2 3 4 - 3
Patient 8 1890 180 0 0 0 1 - 0
@@ V-B Change-point detection for MNIST handwritten digits
To better demonstrate the efficacy of RING-CPD on non-Euclidean datasets, we evaluated its performance using the MNIST dataset of handwritten digits [62]. The MNIST dataset comprises images of digits ranging from 00 to $9$, with each image being a $28\times 28$ matrix. Each matrix element represents a grayscale intensity value between 00 and $255$. Figure 8 presents examples of these images. For our analysis, we select image pairs that are typically challenging to differentiate: digits 00 and $8$, $1$ and $7$, $5$ and $6$, $3$ and $8$, and $4$ and $9$. We structure the dataset such that the first $\tau$ images correspond to one digit and the remaining $n-\tau$ images to another, with configurations $(\tau,n)=(15,30)$ and $(15,45)$. A nominal level $\alpha=0.05$ is used, and each scenario is repeated $100$ times through random sampling of images. The negative Frobenius norm serves as the similarity measure.
We present the empirical power (defined as the proportion of trials with a $p$-value less than $0.05$) in Table VI with the average change-point estimation error $|\tau-\widehat{\tau}|$ for those rejecting $H_{0}$ in parentheses. Notably, the traditional Euclidean distance-based method ED is ineffective across all scenarios, evidenced by a zero empirical detection rate, so we omit its results in Table VI. In contrast, $\rm M_{g}$-NN demonstrates superior performance in nearly all configurations. Other graph-based methodologies like GET, MET, and SWR also perform commendably. However, methods such as DM, $C_{2N}$, and $S_{3}$ perform unsatisfactory in some scenarios, likely due to the sparse and complex nature of image data.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Examples of different digital images.
@@ V-C Changed interval detection for New York taxi data
We illustrate our method for changed interval detection in studying travel pattern changes around New York Central Park. We use the yellow taxi trip records for the year $2014$ for the NYC Taxi and Limousine Commission (TLC) website (https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page), which contains the city’s taxi pickup and drop-off times and locations (longitude and latitude coordinates). We set the latitude range of New York Central Park as $40.76$ to $40.81$ and the longitude range as $-74.10$ to $-73.60$. The boundary of New York City is set as $40.50$ to $40.95$ in latitude and $-74.10$ to $-73.60$. We only consider those trips that began with a pickup in New York City and ended with a drop-off in New York Central Park. We use the two-dimensional kernel density estimation with the bivariate normal kernel and $50$ grid points in each direction to represent the trips of each day in New York City, as illustrated by Figure 9 on two random days. We use the Frobenius norm to construct the similarity graphs in the subsequent analysis. The $p$-values of all methods are obtained through $1000$ random permutations.
@@
We first compare RING-CPD with GET and MET. We set $n_{0}=\max\{5,[0.05n]\}$ and $n_{1}=n-n_{0}$ and the nominal level as $0.01$. All methods detect the same changed interval $06/19$-$08/31$ with $p$-values $<0.001$, which almost overlaps totally with the summer break. Besides, $06/19$ is Juneteenth and $09/01$ is Labor Day.
Since there may be multiple changed intervals, we apply the methods recursively. Specifically, we apply the methods to the three segments divided by the detected changed interval. For the period of $/01$ 01-$06/18$, all methods detect the changed interval $03/18$-$03/28$ with $p$-value $<0.001$, which is around the spring break of most American universities, while $3/17$ is St. Patrick’s Day. For the period of $06/19$-$08/31$, no method can reject the null hypothesis at the $0.01$ significance level. For the period of $09/01$-$12/31$, $\rm M_{g}$-NN, GET, and MET report the changed interval starting at $09/01$, $09/03$, and $09/02$ respectively and all ending at $11/12$ with $p$-values $<0.001$. This is the changed interval of the Fall Semester till Veterans Day (11/$11/1$ 11).
We further apply these methods to the segments divided by themselves longer than $40$ days. The only detected changed interval is $01/27$-$01/31$ reported by $\rm M_{g}$-NN with $p$-value $0.01$ in the segment $/01$ 01-$03/17$, which is around the Lunar New Year ($01/31/2014$). It is worth noting that for the period of $11/13$-$12/31$, all methods report the changed interval that covers Christmas, while $\rm M_{g}$-NN yields a small $p$-value $0.02$ and other methods report $p$-values larger than $0.05$. The results are summarized in Table VII.
@@
$/01$ 01-$12/31$ 06/19-08/31 Summer break
$p$-value $<0.001$
$/01$ 01-$06/18$ 03/18-03/28 Spring break/St. Patrick’s Day
$p$-value $<0.001$
$06/19$-$08/31$ $06/23$-$06/27$ $07/17$-$08/29$ $06/23$-$06/27$
$p$-value $0.048$ $0.196$ $0.223$
$09/01$-$12/31$ 09/01-11/12 09/02-11/12 Fall Semester (till Veterans Day)
$p$-value $<0.001$
$/01$ 01-$03/17$ 01/27-01/31 $02/19$-$03/13$ $02/19$-$03/13$ Lunar New Year
$p$-value $0.01$ $0.051$ $0.263$
$03/29$-$06/18$ $/04$ 04-$06/15$ $04/23$-$05/22$ $04/23$-$05/23$
$p$-value $0.094$ $0.366$ $0.333$
$09/01(03,02)$-$11/12$ $09/02$-$09/11$ $09/03$-$09/11$ $09/02$-$09/12$
$p$-value $0.014$ $0.073$ $0.028$
$11/13$-$12/31$ $12/22$-$12/30$ $12/20$-$12/30$ $12/20$-$12/30$
$p$-value $0.020$ $0.067$ $0.059$
$02/01$-$03/17$ $/03$ 03-$03/07$
$p$-value $0.263$
We also apply other methods to the dataset. Since both ED and $S_{3}$ can detect multiple change points, we apply them directly to the whole sequence $/01$ 01-$12/31$. We also include $C_{2N}$. Although $C_{2N}$ is not designed for multiple change-point detections, we adopt the same binary segmentation procedure $S_{3}$ used [34]. As summarized in Table VIII, ED detects two change-points $06/20$ and $09/02$ both with $p$-values $0.001$. $S_{3}$ only detects one change-point $06/20$ with $p$-value $0.003$. $C_{2N}$ detects two change-points, which are $06/19$ and $09/14$, with $p$-values $<0.001$ and $0.004$. They all miss some important change points shown in Table VII.

### v4

@@ IV-A The choice of k
The choice of graphs remains an open question for CPD based on similarity graphs [chen2015graph, chu2019asymptotic, zhang2021graph]. We adapt the method in [zhang2021graph] and [rise]. Specifically, they compare the empirical power of the method for different choices of $k=[n^{\lambda}]$ by varying $\lambda$ from 0 to $1$. [zhang2021graph] suggested using $k=[n^{0.5}]$ for the generalized edge-count test (GET) [chu2019asymptotic] when the $k$-MST is used, while [rise] recommended $k=[n^{0.65}]$ for their two-sample test statistic with graph-induced rank on the $k$-NNG. We follow the same way in choosing $k$ in $\rm M_{g}$-NN. We generate independent sequences from three different distribution pairs of $(F_{0},F_{1})$:
@@
    Multivariate Gaussian: $\big(N_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),N_{d}(\frac{30}{\sqrt{nd}}\mathbf{1}_{d},\mathbf{I}_{d})\big)$;
@@
    Multivariate $t_{3}$: $\big(Mt_{3}(\mathbf{0}_{d},\mathbf{I}_{d}),Mt_{3}(\frac{30}{\sqrt{nd}}\mathbf{1}_{d},(1+\frac{30}{\sqrt{nd}})^{2}\mathbf{I}_{d})\big)$;
@@
    Multivariate Cauchy: $\big(MC_{d}(\mathbf{0}_{d},\mathbf{I}_{d}),MC_{d}(\frac{30}{\sqrt{n}}\mathbf{1}_{d},\mathbf{I}_{d})\big)$.
The parameters are chosen so that the tests have moderate power. The change-point is set to be $\tau=n/2$, the dimension $d=500$ and $n=50,100,200$. We set $n_{0}=\lceil 0.05n\rceil$ and $n_{1}=n-n_{0}$, which will also be our choice by default in latter experiments, where $\lceil x\rceil$ denotes the smallest integer larger than or equal to $x$. For comparison, we also include two graph-based methods, GET and the max-type edge-count test (MET) proposed in [chu2019asymptotic] using the $k$-MST. The empirical power is defined as the ratio of successful detection where the $p$-value is smaller than $0.05$. For fairness, the $p$-values are approximated by $1,000$ random permutations for all methods.
Figure 6 illustrates the power of the tests for different values of $k=\lfloor n^{\lambda}\rfloor$. For $\rm M_{g}$-NN, $\lambda\in(0,0.95]$, while for GET and MET, $k$ is chosen from $1$ to $\min\{\lfloor n^{0.95}\rfloor,\lfloor\frac{n-1}{2}\rfloor\}$. This range is chosen because the $l$th-MST contains n1n-1 edges, whereas a complete graph has n(n1)/2n(n-1)/2 edges. We observe that the power of these tests initially increases rapidly as $k$ or $\lambda$ increases. However, as $k$ continues to increase, the power of GET and MET decreases sharply. In contrast, $\rm M_{g}$-NN demonstrates greater robustness, maintaining consistent performance even for larger values of $k$ (e.g., $k>n^{0.65}$). In the case of the multivariate Cauchy distribution, the power of $\rm M_{g}$-NN exhibits a slight decrease when $k>n^{0.65}$, but overall remains stable. This robustness of $\rm M_{g}$-NN can be attributed to its use of ranks on the edges, which effectively reduces the impact of additional noise from denser graphs by assigning smaller ranks to edges introduced later. As a result, $\rm M_{g}$-NN consistently achieves the best performance, showing a significant improvement in power for heavy-tailed distributions (such as the multivariate $t_{3}$ and Cauchy distributions) and demonstrating resilience across a wide range of $k$ values. Considering the trade-offs between computational cost, control of type I error, and power, we select $\lambda=0.65$ for $\rm M_{g}$-NN, and $\lambda=0.5$ for GET and MET in subsequent analyses. These choices strike a reasonable balance and are consistent with previous recommendations [zhang2021graph, rise].
@@
Remark 5
Selecting an appropriate value of $k$ remains an open question for graph-based methods in real-world applications. Beyond using a fixed value, such as $k=[n^{0.65}]$, we propose a data-driven approach to empirically determine $k$. Specifically, we first randomly permute the observations, then apply a mean shift to the last $\lceil n/2\rceil$ observations by adding a value $\mu$ to each dimension. We then select the $k$ value that produces the smallest $p$-value. This approach incorporates the unknown underlying distribution for choosing an optimal value of $k$.
@@ IV-B Performance comparison
We compare RING-CPD to GET and MET on $k$-MST with $k=[n^{0.5}]$, the graph-based method on the shortest Hamiltonian path [shi2017consistent] (SWR), the method based on Fréchet means and variances [dubey2020frechet] (DM). We also compare three interpoint distance-based methods, the widely used distance-based method E-Divisive (ED) [matteson2014nonparametric] implemented in the R package ecp, and the other two methods proposed recently by [li2020asymptotic] and [nie2021weighted]. [li2020asymptotic] proposed four statistics and we compare the statistic $C_{2N}$ that had a satisfactory performance in most of their simulation settings. [nie2021weighted] proposed three statistics, which perform well for location change, scale change, and general change, respectively. Here we compare with their statistic $S_{3}$, which they concluded to have relatively robust performance across various alternatives. For fairness, the $p$-values of all these methods are approximated by $1,000$ random permutations.
Alternative
Setting $H_{0}$ (a) (b) (c) (d) (e)
We set $n=200$ and the change-point $\tau=[n/3]$ and consider the dimension of the distributions $d=200,500,1000$. Before the change-point $y_{i}\sim F_{0}$ and after the change-point $y_{i}\sim F_{1}$. We consider both the empirical power and the detection accuracy estimated from $1000$ trials for each scenario. The empirical power is the ratio of the successful detection defined as $p$-value smaller than the nominal level $0.05$. The detection accuracy is provided in parentheses, which is the ratio of trials that the detected change-point is located in $\tau-0.05n,\tau+0.05n]$ and the $p$-value smaller than $0.05$. We consider various settings that cover light-tailed, heavy-tailed, skewed, and mixture distributions for location, scale, and mixed alternatives. Specifically, we consider six settings for $F_{i}$, $i=0,1$:
@@
    Multivariate Gaussian: $N_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})$;
@@
    Multivariate Cauchy: $MC_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})$;
@@
    Multivariate $\chi_{5}^{2}$: $\chi_{5}^{2}(\bm{\mu}_{i},\bm{\Sigma}_{i})$ (generated as $\bm{\Sigma}_{i}^{\frac{1}{2}}(X-5\mathbf{1}_{d}+\bm{\mu}_{i})$ where the $d$ components of $X$ are i.i.d. $\chi_{5}^{2}$);
@@
    Multivariate Gaussian mixture: $WN_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})+(1-W)N_{d}(-\bm{\mu}_{i},\bm{\Sigma}_{i})$ with $W\sim{\rm Bernoulli}(0.5)$;
@@
    Multivariate Gaussian with Multivariate $t_{7}$ outliers $WN_{d}(\bm{\mu}_{i},\bm{\Sigma}_{i})+(1-W)Mt_{7}\big(\bm{\mu}_{i},\bm{\Sigma}_{i})$ with $W\sim{\rm Bernoulli}(0.9)$.
@@ V-A Seizure detection from functional connectivity networks
We illustrate RING-CPD for identifying epileptic seizures, which over two million Americans are suffering from [iasemidis2003epileptic]. As a promising therapy, responsive neurostimulation requires automated algorithms to detect seizures as early as possible. Besides, to identify seizures, physicians have to review abundant electro-encephalogram (EEG) recordings, which in some patients may be quite subtle. Hence, it is important to develop methods with low false positive and false negative rates to detect seizures from the EEG recordings. We use the “Detect seizures in intracranial EEG (iEEG) recordings” database by the UPenn and Mayo Clinic (https://www.kaggle.com/c/seizure-detection), which consists of the EEG recordings of $12$ subjects (eight patients and four dogs). For each subject, both the normal brain activity and the seizure activity are recorded multiple times, which are one-second clips with various channels (from $16$ to $72$), reducing to a multivariate stream of iEEGs. Following the preprocessing procedure of [zambon2019change], we represent the iEEG data as functional connectivity networks using Pearson correlation in the high-gamma band (7070-100100Hz) [bastos2016tutorial]. Functional connectivity networks are weighted graphs, where the vertices are the electrodes, and the weights of edges correspond to the coupling strength of the vertices. An illustration of the networks is in Figure 7. The sample sizes of the $12$ subjects are also different, and the true change-points $\tau$ are also known - before the change-point, the networks are from the seizure period, while after the change-point, the networks are from the normal brain activity, so we have the ground truth. We use the Frobenius norm to measure the distance between the observations represented by the weighted adjacency graphs.
@@
We do not include SWR in the comparison here since SWR does not perform well in the simulation studies and is time-consuming. For $C_{2N}$, since it does well under some simulation settings, we try to include it in the comparison. However, it is not only time-consuming but also memory-consuming (e.g., it requires at least 1717Gb size of memory when $n=1320$); we are only able to run it for $n\leq 600$, thus only showing its result for Dog $1$, and Patients $1$ and $4$. Since the sample size of each subject is large enough, we use the asymptotic $p$-value approximation for $\rm M_{g}$-NN and MET. We omit the result of GET since it performs similarly to MET but its $p$-value approximation is not as exact as MET. For DM, ED, and $S_{3}$, we still use $1,000$ random permutations to obtain the $p$-values. The results are summarized in Table V, where the absolute difference between the true change-point and the detected change-point $|\hat{\tau}-\tau|$ is reported. The $p$-values are not reported as they are smaller than $0.01$ for all methods and subjects. Our method achieves the same detection error as MET, which is very small for all subjects. ED also performs well, but with a slightly large error for Patient $4$. Although DM and $S_{3}$ achieve small errors for most subjects, they attain large detection errors for Patients $3$ and $4$. The performance of $C_{2N}$ is not robust in that it shows a large detection error for Patient $4$.
TABLE V: The absolute difference between the true change-point and the detected change-point ($|\hat{\tau}-\tau|$). The $p$-values of all methods for all subjects (dogs or PTs, where PT== patient) are smaller than $0.01$. $\rm M_{g}$-NN is the proposed method.
$n$ $\tau$ $\rm M_{g}$-NN MET DM ED $C_{2N}$ $S_{3}$
@@
PT 1 174 70 1 1 1 0 7 1
PT 2 3141 151 7 7 13 1 - 13
PT 3 1041 327 0 0 162 1 - 162
PT 4 210 20 0 0 67 11 137 67
PT 5 2745 135 3 3 5 2 - 5
PT 6 2997 225 0 0 0 1 - 0
PT 7 3521 282 2 2 3 4 - 3
PT 8 1890 180 0 0 0 1 - 0
@@ V-B Change-point detection for MNIST handwritten digits
To better demonstrate the efficacy of RING-CPD on non-Euclidean datasets, we evaluated its performance using the MNIST dataset of handwritten digits [lecun1998gradient]. The MNIST dataset comprises images of digits ranging from 0 to $9$, with each image being a $28\times 28$ matrix. Each matrix element represents a grayscale intensity value between 0 and $255$. For our analysis, we select image pairs that are typically challenging to differentiate: digits 0 and $8$, 11 and $7$, 55 and $6$, 33 and $8$, and $4$ and $9$. We structure the dataset such that the first $\tau$ images correspond to one digit and the remaining $n-\tau$ images to another, with configurations $(\tau,n)=(15,30)$ and $(15,45)$. A nominal level $\alpha=0.05$ is used, and each scenario is repeated $100$ times through random sampling of images. The negative Frobenius norm serves as the similarity measure.
[…]
@@ V-C Changed interval detection for New York taxi data
We illustrate our method for changed interval detection in studying travel pattern changes around New York Central Park. We use the yellow taxi trip records for the year $2014$ available on the the NYC Taxi and Limousine Commission (TLC) website (https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page), which contains the city’s taxi pickup and drop-off times and locations (longitude and latitude coordinates). We set the latitude range of New York Central Park as $40.76$ to $40.81$ and the longitude range as 74.10-74.10 to 73.60-73.60. The boundary of New York City is set as $40.50$ to $40.95$ in latitude and 74.10-74.10 to 73.60-73.60. We only consider those trips that began with a pickup in New York City and ended with a drop-off in New York Central Park. We use the two-dimensional kernel density estimation with the bivariate normal kernel and $50$ grid points in each direction to represent the trips of each day in New York City, as illustrated by Figure 8 on two random days. We use the Frobenius norm to construct the similarity graphs in the subsequent analysis. The $p$-values of all methods are obtained through $1000$ random permutations.
@@
01/0101/01-$12/31$ 06/19-08/31 Summer break
$p$-value <0.001<0.001
01/0101/01-$06/18$ 03/18-03/28 Spring break/St. Patrick’s Day
$p$-value <0.001<0.001
06/1906/19-$08/31$ 06/2306/23-$06/27$ 07/1707/17-$08/29$ 06/2306/23-$06/27$ -
pp-value $0.048$ $0.196$ $0.223$
09/0109/01-$12/31$ 09/01-11/12 09/03-11/12 09/02-11/12 Fall Semester (till Veterans Day)
$p$-value <0.001<0.001
01/0101/01-$03/17$ 01/27-01/31 02/1902/19-$03/13$ 02/1902/19-$03/13$ Lunar New Year
pp-value $0.01$ $0.051$ $0.263$
03/2903/29-$06/18$ 04/0404/04-$06/15$ 04/2304/23-$05/22$ 04/2304/23-$05/23$ -
pp-value $0.094$ $0.366$ $0.333$
09/01(03,02)09/01(03,02)-$11/12$ 09/0209/02-$09/11$ 09/0309/03-$09/11$ 09/0209/02-$09/12$ -
pp-value $0.014$ $0.073$ $0.028$
11/1311/13-$12/31$ 12/2212/22-$12/30$ 12/2012/20-$12/30$ 12/2012/20-$12/30$ -
pp-value $0.020$ $0.067$ $0.059$
02/0102/01-$03/17$ 03/0303/03-$03/07$ -
pp-value $0.263$
We first compare RING-CPD with GET and MET. We set $n_{0}=\max\{5,[0.05n]\}$, $n_{1}=n-n_{0}$, and the nominal level to be $0.01$. All methods detect the same changed interval 06/1906/19-$08/31$ with $p$-values <0.001<0.001, which almost entirely overlaps with the summer break. Besides, $06/19$ is Juneteenth and $09/01$ is Labor Day.
Since there may be multiple changed intervals, we apply the methods recursively. Specifically, we apply the methods to the three segments divided by the detected changed interval. For the period of 01/0101/01-$06/18$, all methods detect the changed interval 03/1803/18-$03/28$ with $p$-value <0.001<0.001, which is around the spring break period for most American universities, while $3/17$ is St. Patrick’s Day. For the period of 06/1906/19-$08/31$, no method can reject the null hypothesis at the $0.01$ significance level. For the period of 09/0109/01-$12/31$, $\rm M_{g}$-NN, GET, and MET report the changed interval starting at $09/01$, $09/03$, and $09/02$ respectively and all ending at $11/12$ with $p$-values <0.001<0.001. This is the changed interval of the Fall Semester till Veterans Day ($11/11$).
We further apply these methods to the segments that are longer than $40$ days. The only detected changed interval is 01/2701/27-$01/31$ reported by $\rm M_{g}$-NN with $p$-value $0.01$ in the segment 01/0101/01-$03/17$, which is around the Lunar New Year ($01/31/2014$). It is worth noting that for the period of 11/1311/13-$12/31$, all methods report the changed interval that covers Christmas, while $\rm M_{g}$-NN yields a small $p$-value $0.02$ and other methods report $p$-values larger than $0.05$. The results are summarized in Table VII.
We also apply other methods to the dataset. Since both ED and $S_{3}$ can detect multiple change points, we apply them directly to the whole sequence 01/0101/01-$12/31$. We also include $C_{2N}$. Although $C_{2N}$ is not designed for multiple change-point detections, we adopt the same binary segmentation procedure used by $S_{3}$ [nie2021weighted]. As summarized in Table VIII, ED detects two change-points $06/20$ and $09/02$ both with $p$-values $0.001$. $S_{3}$ only detects one change-point $06/20$ with $p$-value $0.003$. $C_{2N}$ detects two change-points, which are $06/19$ and $09/14$, with $p$-values <0.001<0.001 and $0.004$. They all miss some important change points shown in Table VIII.